# Dynamical stability of evolutionarily stable strategy in asymmetric games

### v2

@@ III.2.2 Non-strict pure NE as ESS
Figure 1: Phase portraits supplementing Examples 1–6 of the main text which should be referred for further details. The red circular marker indicates the ESS (fixed point) under focus. The blue lines are trajectories under replicator equation whereas the green dashed lines are trajectories under adjusted replicator equation.
The case of pure ESS that is non-strict (or weak) NE is laden with subtleties, mainly because the corresponding fixed point happens not to be isolated. For instance, the equivalence between the two definitions (and- and sum-definitions) in the context of linear stability and Lyapunov function, discussed above, may be broken; after all, the linear stability analysis would fail as one eigenvalue should be zero and the Lyapunov function argument would not applicable since the fixed point is not isolated. Below we highlight such points through simple examples.
Example 1 (see Fig. 1a): ESS under sum-definition need not be an ESS under and-definition.
Consider the following general class of bimatrix games:
${\sf B}=\begin{pmatrix}b_{1}&b_{2}\\ 0&b_{3}\end{pmatrix}~{}\textrm{and}~{}{\sf C}=\begin{pmatrix}c_{1}&0\\ c_{1}&c_{3}\end{pmatrix},$ (22)
where $b_{1}>0$ and $c_{3}>0$ such that it has a non-strict pure NE, $(\hat{x},\hat{y})=(1,1)$. To confirm whether it is an ESS, we denote the probability vector corresponding to the NE as $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}1,0\end{pmatrix},\begin{pmatrix}1,0\end{pmatrix}\right)$ and the neighbourhood state as $(\mathbf{x},\mathbf{y})\equiv\left(\begin{pmatrix}1-\epsilon,\epsilon\end{pmatrix},\begin{pmatrix}1-\delta,\delta\end{pmatrix}\right)$. We find that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}=b_{1}\epsilon+\delta\epsilon(-b_{1}+b_{2}-b_{3})$ which is positive for infinitesimally small $\epsilon$ and $\delta$, and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}=-\delta\epsilon c_{3}$ which is negative violating the condition of and-definition. To find whether it satisfies sum-definition of ESS, we calculate: $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}-\mathbf{y}\cdot{\sf C}\mathbf{x}=b_{1}\epsilon+\delta\epsilon(-b_{1}+b_{2}-b_{3}-c_{3})$ which being positive for infinitesimal $\epsilon$ implies $(\hat{x},\hat{y})$ satisfies sum-definition. In passing, one can check that, as has been generally established in Section II.2, it is 2ESS only under simultaneous deviation perspective.
Next, to ascertain the nature of dynamical stability of the ESS, we write the corresponding replicator equation for payoff matrices in (22):
$\dot{x}$ $=$ $x(1-x)\left[b_{2}-b_{3}+y(b_{1}-b_{2}+b_{3})\right],$ (23)
$\dot{y}$ $=$ $-y(1-y)(1-x)c_{3}.$ (24)
We can see that the fixed point $(1,1)$ can not be asymptotically stable because $\dot{y}$—being non-positive in the fixed point’s neighbourhood—stops $y$-value of any trajectory starting in the neighbourhood from approaching $1$. This demonstrates that, in the case of non-strict NE, Definition 1c and 3c (face-value perspective) are better indicator of stability than Definition 2c. This is because failing to satisfy Definition 1c or 3c, renders the fixed point to be unstable while ESS satisfying Definition 2c wrongly predicts the corresponding fixed point to be stable in the light of Theorem 1.
Example 2 (see Fig. 1b): ESS under and-definition need not be an ESS under or-definition if one adopts face-value perspective, but it is an ESS under or-definition with simultaneous deviation perspective.
@@
${\sf B}=\begin{pmatrix}b_{1}&b_{2}\\ b_{1}&b_{3}\end{pmatrix}~{}\textrm{and}~{}{\sf C}=\begin{pmatrix}c_{1}&c_{2}\\ c_{1}&c_{3}\end{pmatrix},$ (25)
where $b_{2}>b_{3}$ and $c_{2}>c_{3}$. We can see that $(\hat{x},\hat{y})=(1,1)$ is a non-strict NE and we can again take neighbourhood state as $(\mathbf{x},\mathbf{y})\equiv\left(\begin{pmatrix}1-\epsilon,\epsilon\end{pmatrix},\begin{pmatrix}1-\delta,\delta\end{pmatrix}\right)$. We find that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}=(b_{2}-b_{3})\epsilon\delta$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}=(c_{2}-c_{3})\epsilon\delta$ which are positive for infinitesimally small $\epsilon$ and $\delta$. Therefore, the NE satisfies and-definition but it does not satisfy or-definition in face-value perspective (since for $\epsilon=0$ both the terms become zero). However, adapting simultaneous deviation perspective puts restrictions that $\epsilon\neq 0$ and $\delta\neq 0$, thereby making it an ESS under or-definition.
@@
It is evident that both $\dot{x}$ and $\dot{y}$ are positive, indicating that the dynamics is driven away from $x=0$ and $y=0$ by any infinitesimal perturbation. Note that since $x=1$ and $y=1$ are line of fixed points, once a trajectory reaches at either boundary, $x=1$ or $y=1$, it terminates. Furthermore, since $\frac{dy}{dx}=\frac{ky}{x}$, where $k\equiv\frac{c_{2}-c_{3}}{b_{2}-b_{3}}$, one arrives at $y=Kx^{k}$ ($K$ being an integration constant) as equation of trajectories. This result shows that only a single trajectory can terminate at $(1,1)$; others terminate at other points on the lines $x=1$ and $y=1$. This implies absence of an open neighbourhood of $(1,1)$ such that all the phase points in the neighbourhood be asymptotically attracted towards $(1,1)$. Concluding, the non-strict NE as ESS need not satisfy Theorem 1.
@@ III.3 Mixed ESS
The state-strategy correspondence—discussed in Section III.1—in the case of mixed states should clearly be such that evolutionarily stable state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ corresponds to evolutionarily stable strategy $(\mathbf{\hat{p}},\mathbf{\hat{q}})$ such that $\mathbf{\hat{p}}=\hat{x}\mathbf{p}_{1}+(1-\hat{x})\mathbf{p}_{2}$ and $\mathbf{\hat{q}}=\hat{y}\mathbf{q}_{1}+(1-\hat{y})\mathbf{q}_{2}$. In this case, as well, a strongly stable $(\mathbf{\hat{p}},\mathbf{\hat{q}})$ would attract $(\sum_{i}x_{i}{\mathbf{p}}_{i},\sum_{i}y_{i}{\mathbf{q}}_{i})$, close to $(\hat{\mathbf{p}},\hat{\mathbf{q}})$. For mixed state case, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$. The former implies
@@
Substituting the matrix elements in terms of ${\sf U}_{1}$ and rearranging the resulting expression we arrive at $\mathbf{\hat{p}}\cdot{\sf U}_{1}\mathbf{\hat{q}}=\mathbf{p}\cdot{\sf U}_{1}\mathbf{\hat{q}}.$ Similarly, $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$ implies $\mathbf{\hat{q}}\cdot{\sf U}_{2}\mathbf{\hat{p}}=\mathbf{q}\cdot{\sf U}_{2}\mathbf{\hat{p}}$. In fact, similar manipulation of terms yield that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}$ implies $\mathbf{\hat{p}}\cdot{\sf U}_{1}\mathbf{q}>\mathbf{p}\cdot{\sf U}_{1}\mathbf{q}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}$ implies $\mathbf{\hat{q}}\cdot{\sf U}_{2}\mathbf{p}>\mathbf{q}\cdot{\sf U}_{2}\mathbf{p}$. Summarizing, if Definition 1b satisfied by the mixed state profile $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, then the conditions of definition are also satisfied by the corresponding strategies $(\mathbf{\hat{p}},\mathbf{\hat{q}})$. Similarly, such correspondence is there for Definitions 2b and 3b as well.
@@
In conclusion, we find that Theorem 1 is invalid if the 2ESS is mixed. The theorem, however, can be saved by either adopting face-value perspective which allows for no mixed ESS or, even more interestingly, using another version of the replicator equation, viz., adjusted replicator equation Mukhopadhyay and Chakraborty (2021); Smith (1982); Hofbauer and Sigmund (1998):
@@
Under its dynamics, the mixed state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ can be asymptotically stable (see Fig. 1c).
Example 3 (see Fig. 1c): Mixed 2ESS under simultaneous deviation perspective is center-type fixed point in $2\times 2$ bimatrix games.
@@
${\sf B}=\begin{pmatrix}\beta_{1}&0\\ 0&\alpha_{1}\end{pmatrix}~{}\text{and}~{}{\sf C}=\begin{pmatrix}\beta_{2}&0\\ 0&\alpha_{2}\end{pmatrix},$ (30)
where for rescaled partnership game, $\alpha_{1}=c\alpha_{2}$ and $\beta_{1}=c\beta_{2}$, on the other hand for rescaled zero-sum game $\alpha_{1}=-c\alpha_{2}$ and $\beta_{1}=-c\beta_{2}$ Hofbauer and Sigmund (1998). Now, the internal fixed point—if it exist in the simplotope—is $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}\alpha_{2}/(\alpha_{2}+\beta_{2}),{\beta_{2}}/({\alpha_{2}+\beta_{2}})\end{pmatrix},\begin{pmatrix}{\alpha_{1}}/({\alpha_{1}+\beta_{1}}),{\beta_{1}}/({\alpha_{1}+\beta_{1}})\end{pmatrix}\right)$ and the neighbourhood vector is $(\mathbf{x},\mathbf{y})=\left(\begin{pmatrix}\alpha_{2}/(\alpha_{2}+\beta_{2})+\delta_{x},{\beta_{2}}/({\alpha_{2}+\beta_{2}})-\delta_{x}\end{pmatrix},\begin{pmatrix}{\alpha_{1}}/({\alpha_{1}+\beta_{1}})+\delta_{y},{\beta_{1}}/({\alpha_{1}+\beta_{1}})-\delta_{y}\end{pmatrix}\right)$, where $\delta_{x}$ and $\delta_{y}$ are infinitesimally small which can have positive or negative sign but neither of them are allowed to be zero as per simultaneous deviation perspective. We find that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{y}=-\delta_{x}\delta_{y}(\alpha_{1}+\beta_{1})$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}=-\delta_{x}\delta_{y}(\alpha_{2}+\beta_{2})$. We can see that the fixed point cannot be an ESS under or-definition if $\alpha_{1}+\beta_{1}$ and $\alpha_{2}+\beta_{2}$ have the same sign, which is the case of rescaled partnership game. The fixed point is always an ESS under or-definition if the they have different sign, this is the case of rescaled zero-sum game. Next we note that the eigenvalues (of Jacobian during linearization about the fixed point) simplify to ±$\pm\sqrt{c}({\alpha_{2}\beta_{2}/(\alpha_{2}+\beta_{2})})$ for rescaled partnership game and ±$\pm i\sqrt{c}({\alpha_{2}\beta_{2}/(\alpha_{2}+\beta_{2})})$ for rescaled zero-sum game, which verifies the claim of ESS being a center.
@@ IV Asymmetric games
It is now natural to ask if introducing some intra-specific interactions changes this situation, i.e., renders Theorem 1—with replicator equation modified to represent the resultant asymmetric game—true for a mixed ESS and makes it possible for a mixed NE to be an ESS under and- and sum-definitions (modified appropriately for asymmetric games). In the present section, we will discuss how adding intra-specific interactions addresses these questions and makes the concept of 2ESS more useful in the context of asymmetric games (see also Appendix A).
Due to the presence of intra-specific interaction, we must introduce two additional square payoff matrices—denoted as $\sf{A}$ and $\sf{D}$, respectively of sizes $m$ and $n$—accounting for the payoff gained by players of population X and players of population Y, respectively. The new payoff components are just the ones that one uses very commonly in symmetric games Smith (1982). Furthermore, rather than working with Definitions 1b, 2b, and 3b, it turns out to be more convenient to work equivalently with Definitions 1c, 2c, and 3c, which however now must be modified to include the intra-specific interactions. Consequently, we write the extensions—Definitions 1d, 2d, and 3d—of Definitions 1c, 2c, and 3c, respectively, as follows:
  1. Definition 1d (and-definition): A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}~{}\text{\emph{and}}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}>\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}.$ (31)
    Taylor Taylor (1979), however, proposes that for $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ to be ESS the total payoff of the resident against the mutant must be higher than that of mutant’s. Mathematically, we may write,
@@
    As discussed in Section II, since the above definitions lead to only pure ESS, Cressman Cressman (1992, 1996) gave a weaker definition of ESS for asymmetric games. It says that, an ESS is a state-pair ($\mathbf{\hat{x}}$,$\mathbf{\hat{y}}$) where, for any small perturbation, at least one of the states is more advantageous. Mathematically,
@@
    $\text{either}~{}~{}\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}~{}~{}\text{\emph{or}}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}>\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}.$ (33)
@@
Let us also, for clarity, reiterate Theorem 1 in the context of asymmetric games Cressman (1992, 1996).
  1. Theorem ${\bm{1}^{\prime}}$: An ESS is a locally asymptotically stable rest point of the following replicator equations,
@@
We now turn to discussing pure and mixed ESS’s separately in the context of asymmetric games.

### v3

@@ III.2.2 Non-strict pure NE as ESS
Figure 1: Phase portraits supplementing Examples 1–5 of the main text which should be referred for further details. The red circular marker indicates the ESS (fixed point) under focus. The blue lines are trajectories under replicator equation whereas the green dashed lines are trajectories under adjusted replicator equation.
The case of pure ESS that is non-strict (or weak) NE is laden with subtleties, mainly because the corresponding fixed point happens not to be isolated. For instance, the equivalence between the two definitions (and- and sum-definitions) in the context of linear stability and Lyapunov function, discussed above, may be broken; after all, the linear stability analysis would fail as one eigenvalue should be zero and the Lyapunov function argument cannot be applicable since the fixed point is not isolated. Below we highlight such points through a simple example.
Example 1 (see Fig. 1a): ESS under and-definition need not be an ESS under or-definition if one adopts face-value perspective, but it is an ESS under or-definition with simultaneous deviation perspective. It is not be an ESS under sum-definition as well.
@@
${\sf B}=\begin{pmatrix}b_{1}&b_{2}\\ b_{1}&b_{3}\end{pmatrix}\leavevmode\nobreak\ \textrm{and}\leavevmode\nobreak\ {\sf C}=\begin{pmatrix}c_{1}&c_{2}\\ c_{1}&c_{3}\end{pmatrix},$ (22)
where $b_{2}>b_{3}$ and $c_{2}>c_{3}$. We can see that $(\hat{x},\hat{y})=(1,1)$ is a non-strict NE and we can again take neighbourhood state as $(\mathbf{x},\mathbf{y})\equiv\left(\begin{pmatrix}1-\epsilon,\epsilon\end{pmatrix},\begin{pmatrix}1-\delta,\delta\end{pmatrix}\right)$. We find that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}=(b_{2}-b_{3})\epsilon\delta$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}=(c_{2}-c_{3})\epsilon\delta$ which are positive for infinitesimally small $\epsilon$ and $\delta$. Therefore, the NE satisfies and-definition but it does not satisfy or-definition in face-value perspective (since for $\epsilon=0$ both the terms become zero). However, adapting simultaneous deviation perspective puts restrictions that $\epsilon\neq 0$ and $\delta\neq 0$, thereby making it an ESS under or-definition. Finally, to justify the non-equivalence of and-and sum-definitions, we calculate $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}-\mathbf{y}\cdot{\sf C}\mathbf{x}=(b_{2}-b_{3})\epsilon\delta+(c_{2}-c_{3})\epsilon\delta$ as one can see that the fixed point $(1,1)$ does not satisfy the sum-definition (for $\epsilon=0$ and $\delta\neq 0$ the term becomes zero).
@@
It is evident that both $\dot{x}$ and $\dot{y}$ are positive, indicating that the dynamics is driven away from $x=0$ and $y=0$ by any infinitesimal perturbation. Note that since $x=1$ and $y=1$ are line of fixed points, once a trajectory reaches at either boundary, $x=1$ or $y=1$, it terminates. Furthermore, since $\frac{dy}{dx}=\frac{ky}{x}$, where $k\equiv\frac{c_{2}-c_{3}}{b_{2}-b_{3}}$, one arrives at $y=Kx^{k}$ ($K$ being an integration constant) as equation of trajectories. This result shows that only a single trajectory can terminate at $(1,1)$; others terminate at other points on the lines $x=1$ and $y=1$. This implies the absence of an open neighbourhood of $(1,1)$ such that all the phase points in the neighbourhood be asymptotically attracted towards $(1,1)$. Consequently, the non-strict NE as ESS need not satisfy Proposition 1.
@@ III.3 Mixed ESS
In the case of mixed states, the state-strategy correspondence—discussed in Section IIIA—shows that evolutionarily stable state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ corresponds to evolutionarily stable strategy $(\mathbf{\hat{p}},\mathbf{\hat{q}})$ such that $\mathbf{\hat{p}}=\hat{x}\mathbf{p}_{1}+(1-\hat{x})\mathbf{p}_{2}$ and $\mathbf{\hat{q}}=\hat{y}\mathbf{q}_{1}+(1-\hat{y})\mathbf{q}_{2}$. In this case, as well, a strongly stable $(\mathbf{\hat{p}},\mathbf{\hat{q}})$ would attract $(\sum_{i}x_{i}{\mathbf{p}}_{i},\sum_{i}y_{i}{\mathbf{q}}_{i})$, close enough to $(\hat{\mathbf{p}},\hat{\mathbf{q}})$. For mixed state case, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$. The former implies
@@
Substituting the matrix elements in terms of ${\sf U}_{1}$ and rearranging the resulting expression we arrive at $\mathbf{\hat{p}}\cdot{\sf U}_{1}\mathbf{\hat{q}}=\mathbf{p}\cdot{\sf U}_{1}\mathbf{\hat{q}}.$ Similarly, $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$ implies $\mathbf{\hat{q}}\cdot{\sf U}_{2}\mathbf{\hat{p}}=\mathbf{q}\cdot{\sf U}_{2}\mathbf{\hat{p}}$. In fact, similar manipulation of terms yield that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}$ implies $\mathbf{\hat{p}}\cdot{\sf U}_{1}\mathbf{q}>\mathbf{p}\cdot{\sf U}_{1}\mathbf{q}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}$ implies $\mathbf{\hat{q}}\cdot{\sf U}_{2}\mathbf{p}>\mathbf{q}\cdot{\sf U}_{2}\mathbf{p}$. Summarizing, if Definition 1b is satisfied by the mixed state profile $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, then the conditions of definition are also satisfied by the corresponding strategies $(\mathbf{\hat{p}},\mathbf{\hat{q}})$. Similarly, there is such correspondence for Definitions 2b and 3b as well.
@@
In conclusion, we find that Proposition 1 is invalid if the 2ESS is mixed. The proposition, however, can be saved by either adopting face-value perspective which allows for no mixed ESS or, even more interestingly, using another version of the replicator equation, viz., adjusted replicator equation Mukhopadhyay and Chakraborty (2021); Smith (1982); Hofbauer and Sigmund (1998):
@@
Under this dynamics, the mixed state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ can be asymptotically stable (see Fig. 1b).
Example 2 (see Fig. 1b): Mixed 2ESS under simultaneous deviation perspective is center-type fixed point in $2\times 2$ bimatrix games.
@@
${\sf B}=\begin{pmatrix}\beta_{1}&0\\ 0&\alpha_{1}\end{pmatrix}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ {\sf C}=\begin{pmatrix}\beta_{2}&0\\ 0&\alpha_{2}\end{pmatrix},$ (27)
where for rescaled partnership game, $\alpha_{1}=c\alpha_{2}$ and $\beta_{1}=c\beta_{2}$, on the other hand for rescaled zero-sum game $\alpha_{1}=-c\alpha_{2}$ and $\beta_{1}=-c\beta_{2}$ Hofbauer and Sigmund (1998). Now, the internal fixed point—if it exist in the simplotope—is $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}\alpha_{2}/(\alpha_{2}+\beta_{2}),{\beta_{2}}/({\alpha_{2}+\beta_{2}})\end{pmatrix},\begin{pmatrix}{\alpha_{1}}/({\alpha_{1}+\beta_{1}}),{\beta_{1}}/({\alpha_{1}+\beta_{1}})\end{pmatrix}\right)$ and the neighbourhood vector is $(\mathbf{x},\mathbf{y})=\left(\begin{pmatrix}\alpha_{2}/(\alpha_{2}+\beta_{2})+\delta_{x},{\beta_{2}}/({\alpha_{2}+\beta_{2}})-\delta_{x}\end{pmatrix},\begin{pmatrix}{\alpha_{1}}/({\alpha_{1}+\beta_{1}})+\delta_{y},{\beta_{1}}/({\alpha_{1}+\beta_{1}})-\delta_{y}\end{pmatrix}\right)$, where $\delta_{x}$ and $\delta_{y}$ are infinitesimally small which can have positive or negative sign but neither of them are allowed to be zero as per simultaneous deviation perspective. We find that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{y}=-\delta_{x}\delta_{y}(\alpha_{1}+\beta_{1})$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}=-\delta_{x}\delta_{y}(\alpha_{2}+\beta_{2})$. We can see that the fixed point cannot be an ESS under or-definition if $\alpha_{1}+\beta_{1}$ and $\alpha_{2}+\beta_{2}$ have the same sign, which is the case of rescaled partnership game. The fixed point is always an ESS under or-definition if they have different sign, this is the case of rescaled zero-sum game. Next we note that the eigenvalues (of Jacobian during linearization about the fixed point) simplify to ±$\pm\sqrt{c}({\alpha_{2}\beta_{2}/(\alpha_{2}+\beta_{2})})$ for rescaled partnership game and ±$\pm i\sqrt{c}({\alpha_{2}\beta_{2}/(\alpha_{2}+\beta_{2})})$ for rescaled zero-sum game, which verifies the claim of ESS being a center.
@@ IV Asymmetric games
It is now natural to ask whether this situation changes introducing some intra-specific interactions, i.e., whether Proposition 1 renders—with replicator equation modified to represent the resultant asymmetric game—true for a mixed ESS and makes it possible for a mixed NE to be an ESS under and- and sum-definitions (modified appropriately for asymmetric games). In the present section, we will discuss how adding intra-specific interactions addresses these questions and makes the concept of 2ESS more useful in the context of asymmetric games.
Due to the presence of intra-specific interaction, we must introduce two additional square payoff matrices—denoted as $\sf{A}$ and $\sf{D}$, respectively of sizes $m$ and $n$—accounting for the payoff gained by players of population X and players of population Y, respectively. The new payoff components are just the ones that one uses very commonly in symmetric games Smith (1982). Furthermore, rather than working with Definitions 1b, 2b, and 3b, it turns out to be more convenient to work with Definitions 1c, 2c, and 3c, which, however, must be modified now to include the intra-specific interactions. Consequently, we write the extensions—Definitions 1d, 2d, and 3d—of Definitions 1c, 2c, and 3c, respectively, as follows:
  1. Definition 1d (and-definition): A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ such that $\mathbf{{x}}\neq\mathbf{\hat{x}}$ and $\mathbf{{y}}\neq\mathbf{\hat{y}}$,
    $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}\leavevmode\nobreak\ \text{\emph{and}}\leavevmode\nobreak\ \mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}>\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}.$ (28)
    Taylor Taylor (1979), however, proposes that for $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ to be an ESS the total payoff of the resident against the mutant must be higher than that of the mutant. Mathematically, we write,
@@
    As discussed in Section II, since the above definitions lead to only pure ESS, Cressman Cressman (1992, 1996) gave a weaker definition of ESS for asymmetric games. According to this definition, an ESS is a state-pair ($\mathbf{\hat{x}}$,$\mathbf{\hat{y}}$) where, for any small perturbation, at least one of the states is more advantageous. Mathematically,
@@
    $\text{either}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{\emph{or}}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}>\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}.$ (30)
@@
Let us also, for clarity, reiterate Proposition 1 refereed as Proposition 1 (cf. Thereom 4 in Ref. Cressman and Tao (2014), Theorem 4.5 and 4.6 in Ref. Cressman (1996) and Theorem 3.4.2 in Ref. Cressman (1992)) in the context of asymmetric games.
  1. Proposition ${\bm{1}^{\prime}}$: An ESS is a locally asymptotically stable rest point of the following replicator equations,
@@
We now turn to discussing pure and mixed ESSes separately in the context of asymmetric games.